A plano-concave lens is made of glass which has a refractive index of 1.5. One surface is flat, and the concave surface has a radius of curvature of 25 cm. determine the focal length of this lens.




1
Expert's answer
2021-07-05T17:48:40-0400

We know that

"\frac{1}{f}=(\mu-1)(\frac{1}{R_2}-\frac{1}{R_2})"

"\frac{1}{f}=(1.5-1)(\frac{1}{infinite}-\frac{1}{25})"

"\frac{1}{f}=0.5(0-\frac{1}{25})"

"\frac{1}{f}=-\frac{1}{50}"

"f=-50 cm"